BSM HQ is back up. I replaced the dovetail with a longer one to help in the balancing, added some cork material to the mounting rings to better hold the telescope tube, and polar aligned the mount.
BSM South and BSM Berry are again operational, as their site managers have returned from their travels.
BSM_NM remains off-line until repair parts for its mount arrive.
OC61 is currently having its mirror realuminized. That process should be completed later this week.
All other systems are operational, weather permitting.